OpenStudy (anonymous):

Square root x+10=x-2

OpenStudy (anonymous):

square both sides of the equation and simplify to get a quadratic equation

OpenStudy (anonymous):

square both sides then factor or use quadratic equation.

OpenStudy (anonymous):

it depends whether the quadratic will factor or not

OpenStudy (anonymous):

x+ 10 = (x-2)^2 x + 10 = x^2 - 4x + 4 x^2 - 5x -6 = 0 will this factor?

OpenStudy (anonymous):

we need two numbers which when multiplied give -6 and when added give -5

OpenStudy (anonymous):

Yes! It factors!! So the quadratic formula and factoring will work. That's so cool!

OpenStudy (anonymous):

yeah!